A body is dropped from a height of 16 m The body strikes the ground and looses 25% of its velocity. The body rebounds to a height of ........

Grade:11

2 Answers

abhi
7 years ago
Let v be the velocity when it stri​kes the ground .ie final velocity ,then v​2-0​2​=2*16*10=320,after rebound velocity V=3/4v,hight attained h=V​2​/2*10=(3/4)​2​*320/20=9/16*16=9m
 
Praveen Mishra
7 years ago
When particle is dropped it will reached at ground with velocity √2gh OK.then it will rebounded with same velocity but of lose of energy it`s velocity becomes 0.75√2gh then after rebounded it`s final velocity and calculated height is 1.5 times of h.